# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.6.1 Average Flow Approximation Series (AFAS)
To simplify notations, we shall write for all $X\in\mathcal{X}(M)$, $\overline{X}[\mu]\coloneqq X-\overline{V}[\mu]$, for all $\mu\in\mathcal{W}_{p}(M)$. For $\mu\in\mathcal{W}_{p}(M)$ and $X\in W^{2,\infty}(M)$, we shall denote the transport plan $(\operatorname{id},\rho^{X}_{\sqrt{t}})_{*}\mu$ just by $\rho^{X}_{\sqrt{t}}$. Then, $\int_{\mathcal{X}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)$ is a transport plan from $\mu$ to $f^{V}_{t}(\mu)$.
@@
$\mathcal{W}_{p}(\mu,f^{V}_{t}(\mu))^{p}$ $\leq\int_{M\times M}d(x,y)^{p}\operatorname{d}\biggl{(}\int_{\mathcal{X}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)\biggr{)}(x,y)$ (3.51)
$=\int_{0}^{\infty}\biggl{(}\int_{\mathcal{X}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)\biggr{)}(\{d(x,y)^{p}>\lambda\})\operatorname{d}\lambda$
$=\int_{0}^{\infty}\int_{\mathcal{X}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}(\{d(x,y)^{p}>\lambda\})\operatorname{d}V[\mu](X)\operatorname{d}\lambda$
$=\int_{\mathcal{X}(M)}\int_{0}^{\infty}\rho^{\overline{X}[\mu]}_{\sqrt{t}}(\{d(x,y)^{p}>\lambda\})\operatorname{d}\lambda\operatorname{d}V[\mu](X)$
$=\int_{\mathcal{X}(M)}\int_{M}d(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x))^{p}\operatorname{d}\mu(x)\operatorname{d}V[\mu](X)$
$=\int_{M}\int_{\mathcal{X}(M)}d(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x))^{p}\operatorname{d}V[\mu](X)\operatorname{d}\mu(x).$
@@
$d(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x))\leq\int_{0}^{\sqrt{t}}|\overline{X}[\mu](\rho^{X}_{t}(x))|_{g}\operatorname{d}t\leq\sqrt{t}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty}).$ (3.52)
@@
$\int_{\mathcal{X}(M)}$ $d(x,\rho^{X}_{\sqrt{t}}(x))^{p}\operatorname{d}V[\mu](X)$ (3.53)
$\leq(\sqrt{t})^{p}\int_{\mathcal{X}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)<\infty,$
@@
$\mathcal{W}_{p}(\mu,f^{V}_{t}(\mu))\leq\sqrt{t}\biggl{(}\int_{\mathcal{X}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr{)}^{\frac{1}{p}}.$ (3.54)
@@
$\mathcal{W}_{p}(f^{V}_{s}(\mu),f^{V}_{t}(\mu))$ $\leq|\sqrt{t}-\sqrt{s}|\biggl{(}\int_{\mathcal{X}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr{)}^{\frac{1}{p}}$ (3.55)
$\leq\sqrt{|t-s|}\biggl{(}\int_{\mathcal{X}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr{)}^{\frac{1}{p}}$
$\leq\sqrt{|t-s|}(\|\overline{V}[\mu]\|_{\infty}+\mathcal{M}_{p}(V[\mu])^{\frac{1}{p}})$
@@
To complete the preparations for our first main result, we enlarge our repertoire of estimates concerning AFAS by some sharper and more technical results than Lemma 3.30. These will allow us to control the speed of divergence of two AFAS with different initial values and different step lengths. However, stronger assumptions on the vector field $V$ and the manifold $(M,g)$ are required for those results.
@@
Assume $(M,g)=(\mathbb{R}^{n},g^{eucl})$. Let $V$ be a vector field probability. Assume that there exist $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]$ is supported in the centered ball with radius $R$. Let $t>0$ and $\mu,\nu,v_{1},v_{2}\in\mathcal{W}_{p}(M)$. Then, if $t\leq 1$, there exists a constant $\mathcal{K}>0$ such that
$\mathcal{W}_{p}(f^{V[v_{1}]}_{t^{2}}(\mu),f^{V[v_{2}]}_{t^{2}}(\nu))^{p}\leq\mathcal{W}_{p}(\mu,\nu)^{p}$ (3.56)
$\qquad+2^{p-3}p\biggl{(}9(p-1)\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}$
[…]
$\qquad+\mathcal{K}\max(1,R^{p+2})\biggl{(}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}$
$\qquad+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])+\mathcal{W}_{p}(\mu,\nu)^{p-2}+\mathcal{W}_{p}(\mu,\nu)^{p-1}\biggr{)}t^{\min(p,3)},$
@@
$\qquad\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-3}3p\mathcal{W}_{p}(\mu,\nu)^{p-1}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t$
$\qquad+\mathcal{K}\max(1,R^{p+1})\biggl{(}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])+\mathcal{W}_{p}(\mu,\nu)^{p-2}$
$\qquad+\mathcal{W}_{p}(\mu,\nu)^{p-1}\biggr{)}t^{\min(p-1,2)}.$
@@
[…]
$\qquad+\mathcal{K}\max(1,R^{p+2})\biggl{(}L^{2}\mathcal{W}_{p}(\mu,\nu)^{2}+L\mathcal{W}_{p}(\mu,\nu)+\mathcal{W}_{p}(\mu,\nu)^{p-2}$
$\qquad+\mathcal{W}_{p}(\mu,\nu)^{p-1}\biggr{)}t^{\min(p,3)},$
@@
$\qquad\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-3}3pL\mathcal{W}_{p}(\mu,\nu)^{p}t$
$\qquad+\mathcal{K}\max(1,R^{p+1})\biggl{(}L\mathcal{W}_{p}(\mu,\nu)+\mathcal{W}_{p}(\mu,\nu)^{p-2}+\mathcal{W}_{p}(\mu,\nu)^{p-1}\biggr{)}t^{\min(p-1,2)}.$
@@
$\qquad+2t^{2}\biggl{(}9\mathcal{W}_{2}(V[v_{1}],V[v_{2}])^{2}+3R\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr{)}$
$\qquad+4R^{2}\biggl{(}R\mathcal{W}_{2}(\mu,\nu)+3\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr{)}\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}\biggr{)}e^{8R^{2}t^{2}}.$
@@
$\mathcal{W}_{2}(f^{V[v_{1}]}_{t^{2}}(\mu),f^{V[v_{2}]}_{t^{2}}(\nu))^{2}\leq\biggl{(}\mathcal{W}_{2}(\mu,\nu)^{2}+2(9L^{2}+3LR)\mathcal{W}_{2}(v_{1},v_{2})^{2}t^{2}$ (3.61)
$\qquad+4R^{2}(R+3L)\mathcal{W}_{2}(\mu,\nu)\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}\biggr{)}e^{8R^{2}t^{2}}.$
@@
$\mathcal{W}_{2}(e^{V[v_{1}]}_{t}(\mu),e^{V[v_{2}]}_{t}(\nu))^{2}\leq\mathcal{W}_{2}(\mu,\nu)^{2}+2\biggl{(}3\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(V[v_{1}],V[v_{2}])$ (3.62)
$\qquad+R\mathcal{W}_{2}(\mu,\nu)^{2}\biggr{)}t$
$\qquad+2R\bigl{(}2R\mathcal{W}_{2}(\mu,\nu)+3\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\bigr{)}t^{2}+\frac{4}{3}t^{3},$
@@
$\mathcal{W}_{2}(e^{V[v_{1}]}_{t}(\mu),e^{V[v_{2}]}_{t}(\nu))^{2}\leq\mathcal{W}_{2}(\mu,\nu)^{2}+2(3L+R)\mathcal{W}_{2}(\mu,\nu)^{2}t$ (3.63)
$\qquad+2R(2R+3L)\mathcal{W}_{2}(\mu,\nu)t^{2}+\frac{4}{3}t^{3}.$
@@
Let $\gamma$ be an optimal transport plan from $\mu$ to $\nu$ and $\Gamma$ be an optimal transport plan from $\tau^{-\overline{V}[v_{1}]}_{*}V[v_{1}]$ to $\tau^{-\overline{V}[v_{2}]}_{*}V[v_{2}]$. First, note that,
$\mathcal{W}_{p}(\tau^{-\overline{V}[v_{1}]}_{*}V[v_{1}],\tau^{-\overline{V}[v_{2}]}_{*}V[v_{2}])\leq\mathcal{W}_{p}(V[v_{1}],V[v_{2}])+\|\overline{V}[v_{1}]-\overline{V}[v_{2}]\|_{W^{2,\infty}(M)}$ (3.64)
$\qquad\leq 3\mathcal{W}_{p}(V[v_{1}],V[v_{2}]),$
by Lemma 3.7. For any two vector fields $X,Y\in\mathcal{X}^{2,\infty}(M)$, denote $\rho^{X,Y}_{t}\coloneqq(\rho^{X}_{t},\rho^{Y}_{t})$. Then, one readily shows that $\bigl{(}\rho^{X,Y}_{t}\bigr{)}_{*}\gamma$ is a transport plan between $\bigl{(}\rho^{X}_{t}\bigr{)}_{*}\mu$ and $\bigl{(}\rho^{Y}_{t}\bigr{)}_{*}\nu$.
@@
$\qquad=\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\biggl{(}\int_{M\times M}\varphi(x,y)\operatorname{d}\bigl{(}\rho^{X,Y}_{t}\bigr{)}_{*}\gamma(x,y)\biggr{)}\operatorname{d}\Gamma(X,Y)$
$\qquad=\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\biggl{(}\int_{M\times M}\varphi(\rho^{X}_{t}(x),\rho^{Y}_{t}(y))\operatorname{d}\gamma(x,y)\biggr{)}\operatorname{d}\Gamma(X,Y),$
for all $\varphi\in\mathcal{C}_{c}(M)$. Then, $\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\bigl{(}\rho^{X,Y}_{t}\bigr{)}_{*}\gamma\operatorname{d}\Gamma(X,Y)$ is a transport plan from $f^{V[v_{1}]}_{t^{2}}(\mu)$ to $f^{V[v_{2}]}_{t^{2}}(\nu)$. First, we assume $(M,g)=(\mathbb{R}^{n},g_{eucl})$. Then, we obtain for all $X,Y\in\mathcal{X}^{2,\infty}(M)$, all $x,y\in M$ and all $s>0$,
@@
$+\operatorname{d}X(\rho^{X}_{s}(x))\biggl{(}\operatorname{d}X(\rho^{X}_{s}(x))\bigl{(}\operatorname{d}X(\rho^{X}_{s}(x))\bigr{)}\biggr{)}\text{ for a.e. $s>0$.}$
Thus,
@@
$\quad=p\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{p-2}\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{>},$
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\bigl{\|}\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\bigr{\|}^{p}\biggr{)}(s)$
$\quad=p(p-2)\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{p-2}\biggl{<}\frac{\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)}{\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\|},\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{>}^{2}$
$\quad+p\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{p-2}\biggl{\|}\frac{\operatorname{d}}{\operatorname{d}t}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{\|}^{2}$
$\quad+p\bigl{\|}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr{\|}^{p-2}\biggl{<}\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr{)}(s)\biggr{>}.$
@@
$\qquad+\int_{0}^{t}(t-s)\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl{(}t\mapsto\bigl{\|}\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\bigr{\|}^{p}\biggr{)}(s)\operatorname{d}s,$

### v2

@@ 3.6.2 Averaged flow operators—Hölder–regularity in time
Let $X\in\mathcal{X}^{2,\infty}(M)$ and $\mu\in\mathcal{W}_{p}(M)$. To simplify notation, we shall write $\overline{X}[\mu]\coloneqq X-\overline{V}[\mu]$, and we shall denote the transport plan $(\operatorname{id},\rho^{X}_{\sqrt{t}})_{*}\mu$ just by $\rho^{X}_{\sqrt{t}}$. Then, $\int_{\mathcal{X}^{2,\infty}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)$ is a transport plan from $\mu$ to $f^{V}_{t}(\mu)$.
@@
$\mathcal{W}_{p}(\mu,f^{V}_{t}(\mu))^{p}$ $\leq\int_{M\times M}d(x,y)^{p}\operatorname{d}\biggl(\int_{\mathcal{X}^{2,\infty}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)\biggr)(x,y)$ (3.50)
$=\int_{0}^{\infty}\biggl(\int_{\mathcal{X}^{2,\infty}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}\operatorname{d}V[\mu](X)\biggr)(\{d(x,y)^{p}>\lambda\})\operatorname{d}\lambda$
$=\int_{0}^{\infty}\int_{\mathcal{X}^{2,\infty}(M)}\rho^{\overline{X}[\mu]}_{\sqrt{t}}(\{d(x,y)^{p}>\lambda\})\operatorname{d}V[\mu](X)\operatorname{d}\lambda$
$=\int_{\mathcal{X}^{2,\infty}(M)}\int_{0}^{\infty}\rho^{\overline{X}[\mu]}_{\sqrt{t}}(\{d(x,y)^{p}>\lambda\})\operatorname{d}\lambda\operatorname{d}V[\mu](X)$
$=\int_{\mathcal{X}^{2,\infty}(M)}\int_{M}d(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x))^{p}\operatorname{d}\mu(x)\operatorname{d}V[\mu](X)$
$=\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}d(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x))^{p}\operatorname{d}V[\mu](X)\operatorname{d}\mu(x).$
@@
$d\bigl(x,\rho^{\overline{X}[\mu]}_{\sqrt{t}}(x)\bigr)\leq\int_{0}^{\sqrt{t}}|\overline{X}[\mu](\rho^{X}_{t}(x))|_{g}\operatorname{d}t\leq\sqrt{t}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty}).$ (3.51)
@@
$\int_{\mathcal{X}^{2,\infty}(M)}$ $d(x,\rho^{X}_{\sqrt{t}}(x))^{p}\operatorname{d}V[\mu](X)$ (3.52)
$\leq(\sqrt{t})^{p}\int_{\mathcal{X}^{2,\infty}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)<\infty,$
@@
$\mathcal{W}_{p}(\mu,f^{V}_{t}(\mu))\leq\sqrt{t}\biggl(\int_{\mathcal{X}^{2,\infty}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr)^{\frac{1}{p}}.$ (3.53)
@@
$\mathcal{W}_{p}(f^{V}_{s}(\mu),f^{V}_{t}(\mu))$ $\leq|\sqrt{t}-\sqrt{s}|\biggl(\int_{\mathcal{X}^{2,\infty}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr)^{\frac{1}{p}}$ (3.54)
$\leq\sqrt{|t-s|}\biggl(\int_{\mathcal{X}^{2,\infty}(M)}(\|\overline{V}[\mu]\|_{\infty}+\|X\|_{\infty})^{p}\operatorname{d}V[\mu](X)\biggr)^{\frac{1}{p}}$
$\leq\sqrt{|t-s|}\bigl(\|\overline{V}[\mu]\|_{\infty}+\mathcal{M}_{p}(V[\mu])^{\frac{1}{p}}\bigr)$
@@
3.6.3 Averaged flow operators—Lipschitz measure–dependence
To complete the preparations for our first main result, we enlarge our repertoire of estimates concerning AFAS by some sharper and more technical results than Lemma 3.32. These will allow us to control the speed of divergence of two AFAS with different initial values and different step lengths. However, stronger assumptions on the vector–field $V$ and the manifold $(M,g)$ are required for those results.
@@ 3.6.3 Averaged flow operators—Lipschitz measure–dependence
Assume $(M,g)=(\mathbb{R}^{n},g^{eucl})$. Let $V$ be a probability vector–field. Assume that there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$, $V[\mu]\in\mathcal{W}_{p+2}(\mathcal{X}^{2,\infty}(M))$ and $\mathcal{M}_{p+2}(V[\mu])\leq R^{p+2}$. Let $t>0$ and $\mu,\nu,v_{1},v_{2}\in\mathcal{W}_{p}(M)$. Then, if $t\leq 1$, there exists a constant $\mathcal{K}>0$ depending only on $p$, such that
$\mathcal{W}_{p}(f^{V[v_{1}]}_{t^{2}}(\mu),f^{V[v_{2}]}_{t^{2}}(\nu))^{p}$ (3.55)
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-2}p\biggl(9(p-1)\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}$
$\hskip 18.49988pt+pR^{2}\mathcal{W}_{p}(\mu,\nu)^{p}+6R\mathcal{W}_{p}(\mu,\nu)^{p-1}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])\biggr)t^{2}$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+2})\biggl(t^{p+2}+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])^{2}t^{p}+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p+1}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)^{p-1}t^{3}+\mathcal{W}_{p}(\mu,\nu)^{p-2}t^{4}+\mathcal{W}_{p}(\mu,\nu)^{2}t^{p}+\mathcal{W}_{p}(\mu,\nu)t^{p+1}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{3}+\mathcal{W}_{p}(\mu,\nu)\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p}\biggr).$
@@
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+p2^{p-2}\biggl(R\mathcal{W}_{p}(\mu,\nu)^{p}+3\mathcal{W}_{p}(\mu,\nu)^{p-1}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])\biggr)t$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+1})\biggl(t^{p+1}+\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p}+\mathcal{W}_{p}(\mu,\nu)^{p-2}t^{3}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)^{p-1}t^{2}+\mathcal{W}_{p}(\mu,\nu)^{2}t^{p-1}+\mathcal{W}_{p}(\mu,\nu)t^{p}$
$\hskip 18.49988pt+\mathcal{W}_{p}(\mu,\nu)\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{p-1}+\mathcal{W}_{p}(\mu,\nu)^{p-2}\mathcal{W}_{p}(V[v_{1}],V[v_{2}])t^{2}\biggr).$
@@
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+2^{p-2}p\biggl(9(p-1)L^{2}+pR^{2}+6RL\biggr)\mathcal{W}_{p}(\mu,\nu)^{p}t^{2}$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+2})\biggl(1+\mathcal{W}_{p}(\mu,\nu)^{p-1}+\mathcal{W}_{p}(\mu,\nu)^{p-2}$
$\hskip 18.49988pt+(L^{2}+L+1)\mathcal{W}_{p}(\mu,\nu)^{2}+(1+L)\mathcal{W}_{p}(\mu,\nu)\biggr)t^{\min(3,p)}.$
@@
$\hskip 18.49988pt\leq\mathcal{W}_{p}(\mu,\nu)^{p}+p2^{p-2}\bigl(R+3L\bigr)\mathcal{W}_{p}(\mu,\nu)^{p}t$
$\hskip 18.49988pt+\mathcal{K}\max(1,R^{p+1})\biggl(1+(1+L)\mathcal{W}_{p}(\mu,\nu)+\mathcal{W}_{p}(\mu,\nu)^{p-2}$
$\hskip 18.49988pt+(1+L)\mathcal{W}_{p}(\mu,\nu)^{p-1}+(1+L)\mathcal{W}_{p}(\mu,\nu)^{2}\biggr)t^{\min(p-1,2)}.$
@@
$\hskip 18.49988pt+2t^{2}\biggl(9\mathcal{W}_{2}(V[v_{1}],V[v_{2}])^{2}+3R\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr)$
$\hskip 18.49988pt+4R^{2}\biggl(R\mathcal{W}_{2}(\mu,\nu)+3\mathcal{W}_{2}(V[v_{1}],V[v_{2}])\biggr)\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}\biggr)e^{8R^{2}t^{2}}.$
@@
$\mathcal{W}_{2}(f^{V}_{t^{2}}(\mu),f^{V}_{t^{2}}(\nu))^{2}\leq\biggl(\mathcal{W}_{2}(\mu,\nu)^{2}+2(9L^{2}+3LR)\mathcal{W}_{2}(\mu,\nu)^{2}t^{2}$ (3.60)
$\hskip 18.49988pt+4R^{2}(R+3L)\mathcal{W}_{2}(\mu,\nu)\frac{t^{3}}{6}+\frac{2R^{4}}{3}t^{4}\biggr)e^{8R^{2}t^{2}}.$
@@
$\mathcal{W}_{2}(e^{V[v_{1}]}_{t}(\mu),e^{V[v_{2}]}_{t}(\nu))^{2}\leq\biggl(\mathcal{W}_{2}(\mu,\nu)^{2}+6\mathcal{W}_{2}(\mu,\nu)\mathcal{W}_{2}(V[v_{1}],V[v_{2}])t$ (3.61)
$\hskip 18.49988pt+6R\mathcal{W}_{2}(V[v_{1}],V[v_{2}])t^{2}\biggr)e^{2Rt},$
@@
$\mathcal{W}_{2}(e^{V}_{t}(\mu),e^{V}_{t}(\nu))^{2}\leq\biggl(\mathcal{W}_{2}(\mu,\nu)^{2}+6L\mathcal{W}_{2}(\mu,\nu)^{2}t+6RL\mathcal{W}_{2}(\mu,\nu)t^{2}\biggr)e^{2Rt}.$ (3.62)
@@
Using Theorem 2.3, let $\gamma$ be an optimal transport plan from $\mu$ to $\nu$ and $\Gamma$ be an optimal transport plan from $\tau^{-\overline{V}[v_{1}]}_{*}V[v_{1}]$ to $\tau^{-\overline{V}[v_{2}]}_{*}V[v_{2}]$. First, note that,
$\mathcal{W}_{p}(\tau^{-\overline{V}[v_{1}]}_{*}V[v_{1}],\tau^{-\overline{V}[v_{2}]}_{*}V[v_{2}])$ (3.63)
$\hskip 18.49988pt\leq\mathcal{W}_{p}(V[v_{1}],V[v_{2}])+\|\overline{V}[v_{1}]-\overline{V}[v_{2}]\|_{W^{2,\infty}(M)}$
$\hskip 18.49988pt\leq 3\mathcal{W}_{p}(V[v_{1}],V[v_{2}]),$
by Lemma 3.9. For any two vector–fields $X,Y\in\mathcal{X}^{2,\infty}(M)$, denote $\rho^{X,Y}_{t}\coloneqq(\rho^{X}_{t},\rho^{Y}_{t})$. Then, one readily shows that $\bigl(\rho^{X,Y}_{t}\bigr)_{*}\gamma$ is a transport plan between $\bigl(\rho^{X}_{t}\bigr)_{*}\mu$ and $\bigl(\rho^{Y}_{t}\bigr)_{*}\nu$.
@@
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\biggl(\int_{M\times M}\varphi(x,y)\operatorname{d}\bigl(\rho^{X,Y}_{t}\bigr)_{*}\gamma(x,y)\biggr)\operatorname{d}\Gamma(X,Y)$
$\hskip 18.49988pt=\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\biggl(\int_{M\times M}\varphi(\rho^{X}_{t}(x),\rho^{Y}_{t}(y))\operatorname{d}\gamma(x,y)\biggr)\operatorname{d}\Gamma(X,Y),$
for all $\varphi\in\mathcal{C}_{c}(M)$. Then, $\int_{\mathcal{X}^{2,\infty}(M)\times\mathcal{X}^{2,\infty}(M)}\bigl(\rho^{X,Y}_{t}\bigr)_{*}\gamma\operatorname{d}\Gamma(X,Y)$ is a transport plan from $f^{V[v_{1}]}_{t^{2}}(\mu)$ to $f^{V[v_{2}]}_{t^{2}}(\nu)$. We obtain for all $X,Y\in\mathcal{X}^{2,\infty}(M)$, all $x,y\in M$ and all $s>0$,
@@
$+\operatorname{d}X(\rho^{X}_{s}(x))\biggl(\operatorname{d}X(\rho^{X}_{s}(x))\bigl(\operatorname{d}X(\rho^{X}_{s}(x))\bigr)\biggr)$
for a.e. $s>0$. Thus,
@@
$\qquad=p\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{p-2}\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr$>,
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\bigl\|\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\bigr\|^{p}\biggr)(s)=p(p-2)\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{p-2}$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\cdot\biggl<\frac{\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)}{\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\|},\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr>^{2}$
$\qquad+p\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{p-2}\biggl\|\frac{\operatorname{d}}{\operatorname{d}t}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr\|^{2}$
$\qquad+p\bigl\|\rho^{X}_{s}(x)-\rho^{Y}_{s}(y)\bigr\|^{p-2}\biggl<\rho^{X}_{s}(x)-\rho^{Y}_{s}(y),\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\biggr)(s)\biggr$>.
@@
$\hskip 18.49988pt+\int_{0}^{t}(t-s)\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}\biggl(t\mapsto\bigl\|\rho^{X}_{t}(x)-\rho^{Y}_{t}(y)\bigr\|^{p}\biggr)(s)\operatorname{d}s.$